OpenStudy (anonymous):

A certain airline requires that the total outside dimensions (length + width + height) of a carry-on bag not exceed 54 inches. Suppose you want to carry on a bag whose length is twice its height. What is the largest volume bag of this shape that you can carry on a flight? (Round your answer to the nearest integer.) in3

OpenStudy (perl):

what did you get

OpenStudy (anonymous):

i have no idea what to do

OpenStudy (perl):

A certain airline requires that the total outside dimensions (length + width + height) of a carry-on bag not exceed 54 inches. Suppose you want to carry on a bag whose length is twice its height. What is the largest volume bag of this shape that you can carry on a flight? (Round your answer to the nearest integer.) in3 I think you can assume the carry on is retangular (a knapsack is hard to determine the dimensions) L +W + H <= 54 V = L*W*H V = 2H*W*H

OpenStudy (perl):

what class is this?

OpenStudy (anonymous):

business calc